An alloy that is 50% gold is to be combined with an alloy that is 75% gold to produce 20 ounces of an alloy that is 60% of gold. How many ounces of each alloy must be used?
Answer by mananth(16949)

percent ---------------- quantity ounces
50 ---------------- x
75 ---------------- 20-x
60.00% ---------------- 20
,,,,
50x+75(20-x)=60* 20
50x+1500 -75x=1200
50x-75x =1200-1500
-25x=-300
/-25
x=12 ounces 50%
8 ounces 75%